Thursday, Mar. 03: DAY Shift

From Hall A Wiki
Jump to: navigation, search

Go back to Daily Run Plan Daily Run Plans


    • Last night, MCC had a trouble by beam loss at 3C00 BLM till around 11PM.
    • After reconstructed optics beam line, we had 10uA for some production runs during OWL.

Specific mission

  • Note: MCC RF recovery (8:30-11:30AM) Down time

1/ Doug: Series of magnet measurement <: 1hr / no beam/ but beam permit
>> Bdl vs. current (decreasing magnet 375A->0A by 25A)
>> HALOG: Report
- No controlled access !
Once RF recovery is done,
2/ BCM calibration (high beam current 30uA(60 ideal) with FC & empty target)
Instruction: DOC1DOC2
- Experts will perform this (Dave/Julie/Thir/...) - Calo. HV off
- If beam is not stable nor acceptable (30uA), we may skip this and go to 3/

Update: 11:30PM/ attempt deliver beam at 1PM
Update: 02:08PM/ MCC has a problem LEM error/IOCIN1 issue, no beam...
Update: 03:00PM/ BCM calibration is postponed (delay beam establishing/Hall-D in/unknown capability of beam current)
Update: 03:10PM/ Start to set beam up with Compton chicane (status?)

3/ Set up beam through Compton chicane
- Dave starts to turn on laser for warming up
4/ Production Data with stable beam current

Upcoming event(Long Term)

  • Beam study (Thursday, 3/10)
  • BCM calibration
  • Target boiling test (60uA)

BPM/Spot++ Check

  • After beam lost for considerable time (like beam study, pass change,...)
  • Ask MCC to run 5uA wth carbon hole target and make sure you can see the hole through Spot++